# G3: An Effective and Adaptive Framework for Worldwide Geolocalization Using Large Multi-Modality Models

### v1

@@ A.3 G3 with Query Image
A.4 Details on Reproduction
We provide codes both in anonymous github  [^4] The link will be released after publication. and supplementary documents. Due to anonymity restrictions and size limitations, we can not contain the checkpoints file (1.6G) and the metadata of the MP16-Pro dataset (around 700MB) in anonymous GitHub or supplementary materials. We will release these data as soon as possible.
A.4.1 File Structure
  • checkpoints: The folder we put the well-trained checkpoints. We have contained the pre-trained checkpoints in supplementary documents.
  • index: The folder we put the index file and retrieval image index file. We have contained the necessary index file for RAG process.
  • data: The folder we put dataset, experimental results, and dataset preprocess scripts.
  • test_files: This folder contains one file to aggregate different generations from multiple prompts.
  • utils: This folder contains model scripts for G3 and GeoCLIP.
  • IndexSearch.py: The script for index building, searching, and evaluation.
  • llm_predict.py: The script for generating LLM predictions (including Img2Loc reproduction).
  • run_G3.py: The codes for Geo-alignment.
  • run_GeoCLIP.py: The codes for GeoCLIP.
A.4.2 Usage Example
G3 contains three modules: Geo-alignment, Geo-diversification, and Geo-verification. Since these modules are completely decoupled, you can reproduce the experimental results at any stage. Here, we will provide examples for each module.
Examples for Geo-alignment: Please run the script python run_G3.py to align image representations with GPS and textual information. This step will output the checkpoint file in the checkpoint directory.
Examples for Geo-diversification: We have uploaded the index file for retrieval images for both the IM2GPS3K and YFCC4K datasets with G3. You can also use IndexSearch.py to generate these files by yourself. Then simply run python llm_predict.py to generate the predictions by LMMs. Remember to set the api_key and base_url in the script to use your personal API service. Finally, run the script test_file/aggregate_llm_predictions.ipynb to aggregate predictions from different prompts.
Examples for Geo-verification: We have also uploaded the generated predictions under the corresponding dataset directories. You can simply run python IndexSearch.py –index=g3_mecarto –dataset=im2gps3k or yfcc4k to perform Geo-verification and evaluate the results.
A.5 Limitations

### v2

@@ A.3 G3 with Query Image
A.4 Limitations
@@ A.5 Details on Baselines
  • PIGEON [5]. PIGEON introduces an innovative approach combining semantic geocell creation, multi-task contrastive pretraining, and a novel loss function, and uniquely enhances guess accuracy through retrieval over location clusters.
A.6 More Experimental Results on Geo-alignment.
In this section, we conduct experiments on incorporating more fine-grained textual descriptions in Geo-alignment. Specifically, in addition to including the city, county, and country information in the textual descriptions of coordinates, we also introduce neighborhood information, which is the most fine-grained data that can be obtained from Nominatim. We use G3-N to denote this variant and keep the other hyperparameters the same as G3. The experimental results on IM2GPS3K are presented in Table 7.
Table 7: Experimental results on textual description granularity on IM2GPS3K.
Methods
Street
1km
City
25km
Region
200km
Country
750km
Continent
2500km
G3-N 16.44 40.64 54.35 70.57 83.98
G3 16.65 40.94 55.56 71.24 84.68
From the results, we can see that G3 outperforms G3-N across all metrics. This may be because the text encoder’s pre-training corpus contains very few instances of neighborhood-level information, resulting in weaker modeling capabilities for neighborhood names. Therefore, introducing neighborhood information into the textual descriptions of coordinates actually adds noise, which negatively impacts the effectiveness of Geo-alignment and subsequently reduces the model’s prediction accuracy.
A.7 Hard Sample and Failure Analysis
Hard Sample. In this section, we give a hard sample to illustrate the effectiveness of G3. Figure 9 shows a man holding an American flag in France, the text on the left says “United States of America” in French. G3 can accurately give the prediction coordinate latitude: 48.8529 and longitude: 2.3632 located in Paris, France. This demonstrates that G3 can effectively avoid the influence of text in images, thereby focusing on the location where the image was taken, showing strong stability.
Figure 9: Hard sample.
Failure Analysis. Figure 10 presents the results of G3 predicting the coordinates of the Eiffel Tower and its replicas. G3 is able to locate the Eiffel Tower in France and its replica in the USA, but fails to distinguish the replica in China. The prediction for the replica in the USA is correct, while the prediction for the replica in China is incorrect. This may be because there are more reference objects around the tower in the USA, which reduces the difficulty of the model’s prediction. In contrast, the lack of reference objects around the tower in China confuses G3’s judgment.
Figure 10: Experimental results of G3 on predicting coordinates of Eiffel Tower and its replica.